x/(x 2 +3x+1)=a (a≠0) find x 2 /(x 4 +3x 2 +1) plzz xplain

x/(x2+3x+1)=a (a≠0)


find x2/(x4+3x2+1)


plzz xplain

Grade:10

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
Please find the answer to your question below
Givenx/(x2+3x+1)=a
So (1/(x+1/x+3))=a
1/a=3+x+1/x
x+1/x=3-(1/a)
squaring on both sides we get
x2+1/x2=7+(1/a2)-(6/a).................(1)
Asked to findx2/(x4+3x2+1)=1/x2+(1/x2)+3
=1/(7-(6/a)+(1/a2)) from (1)
=a2/(7a2-6a+1)